OpenStudy (anonymous):

Consider a radioactive sample with a half-life of one week. How much of the original sample will be left at the end of the second week? The third week? The fourth week? @Abhisar

OpenStudy (abhisar):

Hello @beccamarx19 !

OpenStudy (abhisar):

Half-life (t\(\huge_½\)) is the amount of time required for a quantity to fall to half its value as measured at the beginning of the time period.

OpenStudy (abhisar):

If number of half lives elapsed is n, then fraction of radioactive sample remaining will be \(\huge\frac{1}{2^n}\)

OpenStudy (abhisar):

For instance, if some radioactive substance has a half life of 5 days then after 30 days i.e 6 (n=6 here) half lives \(\huge\frac{1}{2^6}\) i.e \(\huge\frac{1}{64}\) of the original sample will remain

OpenStudy (abhisar):

Can u calculate the answer for ur original question ?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

3.5, 24.5, 10.5?

OpenStudy (abhisar):

In ur question half life is given 1week now for two weeks n=2 Amount left will be \(\huge\frac{1}{2^2}\) = .025 or 25% of the original sample

OpenStudy (abhisar):

Can u calculate for 3rd week ?

OpenStudy (abhisar):

???

OpenStudy (anonymous):

i got 10.5

OpenStudy (anonymous):

3.5,7,10.5

OpenStudy (abhisar):

For 3rd week n will be equal to 3, so amount left = \(\huge\frac{1}{2^3}\) = 0.125 = 12.5%

OpenStudy (anonymous):

i thought it was times the days?

OpenStudy (abhisar):

For 4th week n will be equal to 4, so amount left = \(\huge\frac{1}{2^4}\) = 0.062 = 6.2%

OpenStudy (abhisar):

Times the days was in my example because half life was 5 days, in ur question half life is one week

OpenStudy (anonymous):

hold up

OpenStudy (abhisar):

ok

OpenStudy (anonymous):

ok so it is

OpenStudy (anonymous):

\[1/2^{1}\]

OpenStudy (anonymous):

\[1/2^{2}\] \[1/2^{3}\] ?

OpenStudy (anonymous):

0.5,1,1.5?

OpenStudy (abhisar):

0.5 for 1 week, 0.25 for two weeks, 0.125 for three weeks

OpenStudy (abhisar):

\(\huge\frac{1}{2^2}\) = \(\huge\frac{1}{4}\)

OpenStudy (anonymous):

right

OpenStudy (abhisar):

got it ? u were doing calculationmistake

OpenStudy (anonymous):

and half of 1/4 is 1/8

OpenStudy (anonymous):

GOT IT!!!